1.The government raises taxes by €100 billion. If the marginal propensity to consume is 0.6, what happens to the following? Do they rise or fall? By what amount?


a) Public saving

b) Private saving

c) National saving

d) Investment


1
Expert's answer
2021-05-18T12:52:00-0400

a) Increase in taxes will lead to the increase in public saving by the same amount – 100 billion euro.

b) The increase in taxes will affect disposable income by the same amount – 100 billion euro. With the marginal prosperity to consume 0.6 we can calculate the following: Private saving=100 billion-(0.6-100 billion)= 100 billion-60 billion=40 billion. The private savings fall for 40 billion euro.

c) National saving is sum of both private and public savings, so the national savings will increase for 60 billion euro.

d) National income is described as Y=C(Y-T)+I(r)+G. When we substitute the consumption function and the investment function into national income identity we get :Y-C(Y-T)-G=I(r). From this formula we can see that national savings are equal to the investments. Since national savings increased for 60 billion then investments also increased for the same amount 60 billion.